OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, 7) with slope –2.

OpenStudy (anonymous):

A y – 7 = –2(x – 6) B y = –2x + 7 C y = –2x + 19 D y – 6 = –2(x – 7)

OpenStudy (anonymous):

Use the point-slope form \[y-y_1=m(x-x_1)\] where \[(x_1,y_1)\] is a point on the line and m=the slope of the line